OpenStudy (anonymous):

.Jim opened an account with $400. The account pays three percent quarterly. How much is in the account at the end of two years? Round to the nearest dollar

OpenStudy (lgbasallote):

i believe you apply the formula I = Prt

OpenStudy (anonymous):

Really ?? @lgbasallote

OpenStudy (valpey):

There is a bit of a semantic bug here when you say "three percent quarterly". It could mean 3% is added each quarter or it could be 3% annual interest compounded quarterly.

OpenStudy (anonymous):

two years is 8 quarters Assuming simple interest, A = P(1+rt) = 400(1+8*.03) = 496

OpenStudy (lgbasallote):

idk..like i said i believe lol

OpenStudy (lgbasallote):

i didnt say it was ;) hehe

OpenStudy (valpey):

For the first case we have: \[A=P(1+r)^{nt}\] For the second case we have: \[A=P(1+\frac{r}{n})^{nt}\]

OpenStudy (anonymous):

its the second one.

OpenStudy (valpey):

In both of these examples, t is the number of years and n is the number of periods per year. P is the beginning principal, and A is the Accumulated total. r is the interest rate.

OpenStudy (anonymous):

it is a good question 3% each quarter, or a rate of 3% compounded quarterly the first is a more accurate reading of the words the second is more accurate reading of real life

OpenStudy (valpey):

Most textbooks are careful to differentiate, but some question writers get sloppy.

OpenStudy (valpey):

I would assume the second one.